[Brahms-l] Shift report 20040223 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 23 2004 - 23:49:24 EST
------------------------------------------------------------
Date: Mon Feb 23 23:49:24 2004, Shift: 16:00-24:00
Supervisor: BHS, with: BSN
Log book#15 pages 66 - 66

Shift Summary:
------------------------------------------------------------
1: Machine has been very good indeed - one store in progress when we came on, dumped at 17.30, then another one from 18.00 to 22.40, and now one more from 23.10 still in progress.

2: Data runs at 4deg 1/1A, 40deg 6A (10557-10563, 10564 in progress) - 600k tr.2, 900k tr. 3. 

3: No problems :-)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 23 23:50:12 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 23 2004 - 23:50:35 EST